OpenStudy (anonymous):

Rationalize the denominator of √50 ---- √5 A. 2√5(over)5 B. 5√5(over)2 C. √10 D. 10

OpenStudy (anonymous):

You don't even have to rationalize it...\[\frac{\sqrt{50}}{\sqrt{5}} \implies \sqrt{\frac{50}{5}} \implies?\]

OpenStudy (anonymous):

?not following srry im rllly bad in math:/

OpenStudy (anonymous):

What is 50 ÷ 5 and just put it under the √ symbol.

OpenStudy (anonymous):

√10

OpenStudy (anonymous):

Yup...